# Crossover of Critical Behavior in Dynamic Phase Transitions of Ferromagnetic Thin Films

### v1

@@ Abstract
We investigate the crossover of critical behavior for the dynamic phase transition (DPT) in ferromagnetic thin films using Monte Carlo simulations of the kinetic Ising model, focusing on the scaling behavior of the dynamic order parameter under a time-dependent external magnetic field. Specifically, we study the transition of the critical behavior of such film systems from two-dimensional (2D) to three-dimensional (3D) as a function of the film thickness and the distance to the critical point, which enables dimensional crossover observations. Our results indicate that the effective critical exponents exhibit a clear transition in their scaling behavior, with thinner films showing 2D-like characteristics and thicker films displaying 3D-like behavior, for both the DPT and the thermodynamic phase transitions (TPT). Quantitatively, the crossover from 2D to 3D behavior occurs at larger film thicknesses for the DPT compared to the TPT, suggesting that DPT and TPT are governed by distinctly different length scales and underlying surface effects. These findings are in agreement with experimental observations in ultrathin Co films, where dynamic and thermodynamic critical exponents were found to differ. Therefore, our study provides an in-depth explanation for critical phenomena in thin-film ferromagnets driven by a time-dependent magnetic field. By comparing the dimensional crossover properties of both TPT and DPT, we present a comprehensive understanding of how thin-film geometry and surface effects influence the scaling laws and critical behavior in nonequilibrium systems.
@@ I Introduction
To explain this intriguing result and seeming contradiction, the authors of [34] noted that their films were several monolayers thick and thus represented a thickness regime, in which dimensional crossover from 2D to 3D type behavior had been observed for the TPT, both theoretically and experimentally [35]. So, if TPT and DPT exhibit a difference in the thickness range, at which this dimensional crossover takes place, the seemingly contradictory experimental results could be explained. This implies that upon using the same relative critical regime criteria, the characteristic length-scale, at which the critical exponents change from the 2D to the 3D case would be different in the TPT and DPT. Accordingly, a single film of finite thickness could exhibit different exponents for both types of phase transitions.
@@ II Simulation framework
We simulate the ferromagnetic thin film system described by the Hamiltonian in Eq. (1) where the external magnetic field has a square-wave profile with an amplitude $H_{0}$ and a period $P$. In our Monte Carlo simulations, the boundary conditions are chosen to reflect the physical constraints of real thin-film systems, as shown schematically in Fig. (1). Specifically, periodic boundary conditions are imposed in the $x$ and $y$ directions to mimic a laterally infinite film system and minimize finite-size effects. Free boundary conditions are employed along the $z$ direction, allowing for the consideration of surface effects that are essential in modeling the thin-film geometry. The system we study consists of $N=L_{x}\times L_{y}\times L_{z}$ spins, where $L_{x}$ and $L_{y}$ denote the linear sizes of the system in the $x$ and $y$ directions while $L_{z}$ represents the thickness of the film along the $z$ direction. During the simulations, the length of the $x$ edge of the thin films will be chosen equal to the length of the $y$ edge, and they will be represented as $L_{x}=L_{y}=L$. Such kind of geometry is convenient for modeling thin films of varying thicknesses and lateral dimensions and makes it possible to analyze size-dependent properties, surface effects, and dimensionality-driven phenomena. We note that our data evaluations are conducted across 100 independent simulation runs, and to compute error bars, we use the Jackknife method [41] for both TPT and DPT. Since we aim to compare the dimensional crossover phenomena appearing in TPT and DPT, the simulation frameworks for these two different cases will be presented in separate subsections.
@@ II.1 Simulation protocols and extracted quantities for TPT analysis
[…]
@@
$\beta_{eff}=\frac{\partial\;\text{Log}M}{\partial\;\text{Log}\varepsilon}\Bigr{\rvert}_{H\rightarrow 0},$ (9)
where $\varepsilon$ denotes the temperature deviation relative to $T_{C}$ for each $L_{z}$ value. Specifically, $\varepsilon=[T_{C}(L_{z})-T]/T_{C}(L_{z})$, and $T$ is the temperature of the system. The corresponding results are shown in Fig. 2(b) and will be discussed in Sec. III.
@@ II.2 Simulation protocols and extracted quantities for DPT analysis
To study the dynamic phase transitions, we carry out Monte Carlo simulations on the thin film system using the single-site update Metropolis algorithm [41, 49]. To enhance the efficiency of the simulations, we implement a geometric parallelization scheme, in which the lattice is divided into multiple strips. The division is based on both the thickness of the system and the number of available processors such that the computational load is distributed optimally across the parallelized system. The protocol during the simulations is as follows: We discard the first $10^{3}$ periods of the external field to allow the system to thermalize. Numerical data are then collected and analyzed during the subsequent $10^{4}$ periods of the external field cycle. As mentioned above, a square-type magnetic field has been selected as the source of the time-dependent magnetic field due to its lower computational cost [38]. During the simulations, the times and, consequently, the period of the external magnetic field are expressed in terms of MC steps. The time, at which the system’s initial configuration is prepared is selected as the initial time. When the Metropolis algorithm is applied to one sweep of the lattice, an MC step is performed during which the external magnetic field is fixed. Specifically, the field value is held constant for $P_{1/2}$ MC steps, after which it is inverted. Here, $P_{1/2}$ is the half-period of the external magnetic field. In the new MC step, all spins in the system experience the corresponding value of the external magnetic field at that moment, and this process continues in the same manner. Therefore, as the simulation progresses, the system experiences the effect of an external magnetic field that varies periodically with an amplitude between the values of $H_{0}$ and $-H_{0}$.
@@
$\tilde{\beta}_{eff}=\frac{\partial\;\text{Log}Q}{\partial\;\text{Log}\tilde{\varepsilon}}\Bigr{\rvert}_{H_{b}\rightarrow 0}$ (11)
where $Q$ is the dynamic order parameter defined in Eq. (2), and $\tilde{\varepsilon}$ denotes the half-period deviation relative to $P_{1/2}^{c}$ for each $L_{z}$ value, i.e., $\tilde{\varepsilon}=[P_{1/2}^{c}(L_{z})-P_{1/2}]/P_{1/2}^{c}(L_{z})$. Here, $P_{1/2}^{c}(L_{z})$ is the critical half-period of the system.
@@ III Results and discussion
Figure 3: Effective critical exponent $\tilde{\beta}_{eff}$ in the vicinity of the dynamic phase transition point, calculated from Eq. (11) for varying values of the film thickness from $L_{z}=2$ to $20$. The critical exponents for both two-dimensional and three-dimensional systems are indicated by red dashed lines.
@@
In Fig. 3, the variation of $\tilde{\beta}_{eff}$ is shown as a function of $\tilde{\varepsilon}$ for the DPT in the same film thickness range as for the TPT. In this case, the control parameter is the half-period of the external magnetic field. Figures 2(b) and 3 display fundamentally and qualitatively similar behavior. Specifically, they illustrate a transition from 3D to 2D behavior that depends on the thickness of the system as well as parameters $\varepsilon$ and $\tilde{\varepsilon}$, i.e. the respective distances to the critical point. It is, however, evident from a comparison of Fig. 3 with Fig. 2(b) that the dimensional transition behavior from 3D to 2D criticality is not universal, if one considers the quantitative evolution of the critical exponents $\beta_{eff}$ and $\tilde{\beta}_{eff}$. For instance, the $L_{z}=10$ curves in the DPT stay very close to the pure 2D behavior, whereas in the TPT case, this system comes already quite close to 3D type behavior at intermediate distances to the critical point. Thus, quantitatively, the dimensional crossover exhibits significantly different behavior in TPT and DPT. To illustrate this clearly, we plot the maximum values of the effective critical exponents ($\beta_{eff}^{max}$, $\tilde{\beta}_{eff}^{max}$) as a function of the thickness $L_{z}$ of the system for both the TPT and DPT in Fig. 4(a). As evident from the figure, the $\beta_{eff}^{max}$ and $\tilde{\beta}_{eff}^{max}$ values are strongly dependent on $L_{z}$. The fundamental similarity between TPT and DPT can be seen in the fact that for $L_{z}=2$ and $L_{z}=20$, the numerical values are indeed very similar. However, for intermediate thicknesses, there is a rather significant deviation in between both, mainly showing that TPT starts to approach 3D behavior already for much thinner films, whereas the DPT preserves almost pure 2D behavior until about $L_{z}=10$. This then also explains very well the prior experimental observations of finding 2D DPT behavior and near 3D TPT behavior in the very same thin film sample [34].
@@
To further elucidate the origin of these quantitative differences in between TPT and DPT in terms of their dimensional crossover behavior, the effective critical exponents $\beta_{eff}$ and $\tilde{\beta}_{eff}$ are displayed as a function of the magnetization $M$ and the dynamic order parameter $Q$ in Figs. 5 (a)-(b) for various values of thickness $L_{z}$ of the film. When comparing these two figures, we observe not only that the behavior is not only fundamentally and qualitatively similar, but Fig. 5 also illustrates that the dimensional crossover takes place at fairly comparable values of the order parameter when comparing the TPT and DPT. This observation is particularly intriguing given the very significant differences in the reduced distances to the critical point, represented as $\varepsilon^{*}$ and $\tilde{\varepsilon}^{*}$ in Fig. 4 (b). This suggests that the dimensional crossover point for each particular film, identified by $\beta_{eff}^{max}$ and $\tilde{\beta}_{eff}^{max}$, is rather closely correlated to reaching an intermediate level of the respective order parameter, namely about 60 - 80 per cent of its saturation value. Apparently, these values are achieved already at a much greater relative distance to the critical point in the dynamic case, which then explains the reduced ability of DPT systems to exhibit 3D scaling for intermediate films thicknesses $L_{z}$. It is very likely that this difference in quantitative behavior is associated with relevantly different surface behaviors in both cases, which is consistent with earlier findings that demonstrated that the surface behavior in the DPT can deviate significantly from that of the TPT [37, 18].
@@ IV Conclusions
To conclude, this study presents a detailed investigation into the critical behavior and dimensional crossover of DPT in ferromagnetic thin films, using Monte Carlo simulations of the kinetic Ising model. By systematically analyzing the scaling behavior of the dynamic order parameter and comparing it with TPT, we have illuminated the underlying mechanisms that differentiate these two types of transitions.
Our results reveal that while DPT and TPT share commonalities in their scaling laws and critical exponents, the dimensional crossover for the DPT occurs at significantly larger thicknesses than for the TPT. This distinction arises from the relevantly different length-scale and governing principles of the two transitions, with DPT being influenced by time-dependent external fields and non-equilibrium dynamics, whereas TPT adheres to equilibrium-driven processes. Our findings are furthermore consistent with experimental observations in ultrathin Co films, where the critical exponents for the TPT align with three-dimensional behavior, while those for the DPT retain two-dimensional characteristics [34].
Beyond advancing our understanding of critical phenomena, this work underscores the importance of exploring dimensional crossover in non-equilibrium systems. We believe that this work lays the groundwork for future studies exploring the impact of surface and finite-size effects on the scaling behavior of both thermodynamic and dynamic phase transitions in low-dimensional systems. Further investigations into the detailed mechanisms underlying dimensional crossover for DPT and TPT, especially the role of surfaces in thin films, will yield an even deeper understanding of the nature of these phase transitions in out-of-equilibrium systems.

### v2

# Crossover of Critical Behavior in Dynamic Phase Transitions of Multilayer Ising Model Systems

@@ Abstract
We investigate the crossover of critical behavior for the dynamic phase transition (DPT) in ferromagnetic thin films using Monte Carlo simulations of the kinetic Ising model, focusing on the scaling behavior of the dynamic order parameter under a time-dependent external magnetic field. Specifically, we study the transition of the critical behavior of such multilayer film systems from two-dimensional (2D) to three-dimensional (3D) as a function of the film thickness and the distance to the critical point, which enables dimensional crossover observations. Our results indicate that the effective critical exponents exhibit a clear transition in their scaling behavior, with thinner films showing 2D-like characteristics and thicker films displaying 3D-like behavior, for both the DPT and the thermodynamic phase transitions (TPT). Quantitatively, the crossover from 2D to 3D behavior occurs at larger film thicknesses for the DPT compared to the TPT, suggesting that DPT and TPT are governed by distinctly different length scales and underlying surface effects. These findings are in agreement with experimental observations in ultrathin Co films, where dynamic and thermodynamic critical exponents were found to differ. Therefore, our study provides an in-depth explanation for critical phenomena in thin-film ferromagnets driven by a time-dependent magnetic field. By comparing the dimensional crossover properties of both TPT and DPT, we present a comprehensive understanding of how thin-film geometry and surface effects influence the scaling laws and critical behavior in nonequilibrium systems.
@@ I Introduction
To explain this intriguing result and seeming contradiction, the authors of [34] noted that their films were several monolayers thick and thus represented a multilayer thickness regime, in which dimensional crossover from 2D to 3D type behavior had been observed for the TPT, both theoretically and experimentally [35]. So, if TPT and DPT exhibit a difference in the thickness range, at which this dimensional crossover takes place, the seemingly contradictory experimental results could be explained. This implies that upon using the same relative critical regime criteria, the characteristic length-scale, at which the critical exponents change from the 2D to the 3D case would be different in the TPT and DPT. Accordingly, a single film of finite thickness could exhibit different exponents for both types of phase transitions.
@@ II Simulation framework
We simulate the ferromagnetic thin film system described by the Hamiltonian in Eq. (1) where the external magnetic field has a square-wave profile with an amplitude $H_{0}$ and a period $P$. In our Monte Carlo simulations, the boundary conditions are chosen to reflect the physical constraints of real thin-film systems, as shown schematically in Fig. (1). Specifically, periodic boundary conditions are imposed in the $x$ and $y$ directions to mimic a laterally infinite film system and minimize finite-size effects. Free boundary conditions are employed along the $z$ direction, allowing for the consideration of surface effects that are essential in modeling the thin-film geometry. The system we study consists of $N=L_{x}\times L_{y}\times L_{z}$ spins, where $L_{x}$ and $L_{y}$ denote the linear sizes of the system in the $x$ and $y$ directions while $L_{z}$ represents the number of layers and thus the thickness of the multilayer film along the $z$ direction. During the simulations, the length of the $x$ edge of the thin films will be chosen equal to the length of the $y$ edge, and they will be represented as $L_{x}=L_{y}=L$. Such kind of geometry is convenient for modeling thin films of varying thicknesses and lateral dimensions and makes it possible to analyze size-dependent properties, surface effects, and dimensionality-driven phenomena. We note that our data evaluations are conducted across 100 independent simulation runs, and to compute error bars, we use the Jackknife method [41] for both TPT and DPT. Since we aim to compare the dimensional crossover phenomena appearing in TPT and DPT, the simulation frameworks for these two different cases will be presented in separate subsections.
@@ II.1 Simulation protocols and extracted quantities for TPT analysis
Since the magnetization can take both positive and negative values in the absence of an external magnetic field, we evaluate the absolute value of $M$ when calculating the order parameter.
@@
where $\langle\cdots\rangle$ denotes the thermal average. Each simulation consists of $5\times 10^{4}$ Monte Carlo steps (MCSs) per temperature, providing sufficient data for a meaningful analysis. To guarantee that the system reaches equilibrium before data are collected, the first $10^{4}$ of the MCSs are discarded to facilitate the thermalization process. Simulations are conducted for system sizes of $L=256$ and $L=512$, with calculations performed for various values of the film thickness $L_{z}$ which allows for a detailed examination of the interplay between system size and critical behavior. Numerical results for the Curie temperature $T_{C}$ as a function of $L_{z}$ are depicted in Fig. 2(a). Here, the Curie temperatures for 2D and 3D Ising models with periodic boundary conditions in all directions are denoted by red dashed lines [44, 45]. It is clear from the figure, that upon increasing the thickness $L_{z}$ starting from a 2D monolayer $(L_{z}=1)$, the Curie temperature rises steeply driven by the enhanced stability of the system due to effectively stronger interactions and an accordingly reduced impact of thermal fluctuations. We can also see that even for thicknesses $L_{z}$ of less than 10 layers, the Curie temperature approaches the 3D Ising case already, which implies that also the critical behavior and exponents will be relevantly impacted, even if a finite thickness film is formally a 2D system. At this point, we note that the numerical outcomes reported here for the Curie temperatures are in good agreement with previous studies [46, 47, 48].
@@
$\beta_{eff}=\frac{\partial\;\text{Log}M}{\partial\;\text{Log}\varepsilon}\Bigr{\rvert}_{H=0},$ (9)
where $\varepsilon$ denotes the temperature deviation relative to $T_{C}$ for each $L_{z}$ value that can be represented by the following equation:
$\varepsilon=\frac{T_{C}(L_{z})-T}{T_{C}(L_{z})}.$ (10)
Here, $T$ is the temperature of the system. The corresponding results are shown in Fig. 2(b) and will be discussed in Sec. III.
@@ II.2 Simulation protocols and extracted quantities for DPT analysis
To study the dynamic phase transitions, we carry out Monte Carlo simulations on the thin film system using the single-site update Metropolis algorithm [41, 49]. To enhance the efficiency of the simulations, we implement a geometric parallelization scheme, in which the lattice is divided into multiple strips. The division is based on both the thickness of the system and the number of available processors such that the computational load is distributed optimally across the parallelized system. The protocol during the simulations is as follows: We discard the first $10^{3}$ periods of the external field to allow the system to thermalize. Numerical data are then collected and analyzed during the subsequent $10^{4}$ periods of the external field cycle. As mentioned above, a square-type magnetic field has been selected as the source of the time-dependent magnetic field due to its lower computational cost [38]. During the simulations, the times and, consequently, the period of the external magnetic field are expressed in terms of MC steps. The time, at which the system’s initial configuration is prepared is selected as the initial time. When the Metropolis algorithm is applied to one sweep of the lattice, an MC step is performed during which the external magnetic field is fixed. Specifically, the field value is held constant for $P_{1/2}$ MC steps, after which it is inverted. Here, $P_{1/2}$ is the half-period of the external magnetic field. In the new MC step, all spins in the system experience the corresponding value of the external magnetic field at that moment, and this process continues in the same manner. Therefore, as the simulation progresses, the system experiences the effect of an external magnetic field that varies periodically with an amplitude between the values of $H_{0}$ and $H_{0}$.
@@
$\tilde{\beta}_{eff}=\frac{\partial\;\text{Log}Q}{\partial\;\text{Log}\tilde{\varepsilon}}\Bigr{\rvert}_{H_{b}=0}$ (12)
where $Q$ is the dynamic order parameter defined in Eq. (2), and as in the thermodynamic case, the dynamic order parameter $Q$ is evaluated using its absolute value to properly account for the symmetry of the system under a periodically changing magnetic field. $\tilde{\varepsilon}$ denotes the half-period deviation relative to $P_{1/2}^{c}$ for each $L_{z}$ value, and it can be defined as follows:
$\tilde{\varepsilon}=\frac{P_{1/2}^{c}(L_{z})-P_{1/2}}{P_{1/2}^{c}(L_{z})},$ (13)
with $P_{1/2}^{c}(L_{z})$ being the critical half-period of the system. We note that although $\varepsilon$ and $\tilde{\varepsilon}$ describe different physical quantities (temperature and half-period, respectively), they both quantify the relative distance from the critical point using the natural control parameter of their respective transitions. Hence, their functional roles are conceptually analogous and allow for meaningful comparison when analyzing the critical behavior and crossover properties of TPT and DPT systems.
As described in Eqs. (9) and (12), we employ the direct fitting method to estimate effective critical exponents over a finite temperature (half-period) range near the transition point. While it is well known that this method has limitations in accurately determining asymptotic critical behavior, it remains a widely used approach in the analysis of crossover phenomena, where the focus is on the evolution of effective exponents. As discussed in Refs. [47, 48], the direct method offers a straightforward way to capture the temperature-dependent behavior of the order parameter and to identify trends associated with changes in universality class or transition type.
@@ III Results and discussion
At this point, we also note that the use of effective critical exponents $\beta_{eff}$, as defined in Eq. (9), is essential for identifying the $\varepsilon$-range where power-law scaling of the order parameter is valid. While the value at $\varepsilon=0$ coincides with the actual critical exponent, this value alone is insufficient to characterize the critical region unless the scaling behavior persists across a finite region adjacent to the critical point. In this regard, $\beta_{eff}$, serves as a diagnostic tool for detecting the limits of asymptotic scaling. A constant or nearly constant value of $\beta_{eff}$, over a finite $\varepsilon$-interval, implies the presence of a well-defined critical region. Conversely, strong variations in $\beta_{eff}$ reflect deviations from this regime, typically associated with crossover effects. As seen in Fig. 2(b), the $L_{z}=2$ case displays a broad region with consistent exponent values, validating its effective two-dimensional nature. For larger $L_{z}$, however, the effective exponent varies significantly with $\varepsilon$, indicating the onset of dimensional crossover. This variation provides a clear quantitative indicator of the crossover phenomenon that is occurring in sufficiently thick films.
@@
Figure 3: Effective critical exponent $\tilde{\beta}_{eff}$ in the vicinity of the dynamic phase transition point, calculated from Eq. (12) for varying values of the film thickness from $L_{z}=2$ to 2020. The critical exponents for both two-dimensional and three-dimensional systems are indicated by red dashed lines.
@@
In Fig. 3, the variation of $\tilde{\beta}_{eff}$ is shown as a function of $\tilde{\varepsilon}$ for the DPT in the same film thickness range as for the TPT. In this case, the control parameter is the half-period of the external magnetic field. Figures 2(b) and 3 display fundamentally and qualitatively similar behavior. Specifically, they illustrate a transition from 3D to 2D behavior that depends on the thickness of the system as well as parameters $\varepsilon$ and $\tilde{\varepsilon}$, i.e. the respective distances to the critical point. It is, however, evident from a comparison of Fig. 3 with Fig. 2(b) that the dimensional transition behavior from 3D to 2D criticality is not universal, if one considers the quantitative evolution of the critical exponents $\beta_{eff}$ and $\tilde{\beta}_{eff}$. For instance, the $L_{z}=10$ curves in the DPT stay very close to the pure 2D behavior, whereas in the TPT case, this system comes already quite close to 3D type behavior at intermediate distances to the critical point. Thus, quantitatively, the dimensional crossover exhibits significantly different behavior in TPT and DPT. To illustrate this clearly, we plot the maximum values of the effective critical exponents ($\beta_{eff}^{max}$, $\tilde{\beta}_{eff}^{max}$) as a function of the thickness $L_{z}$ of the system for both the TPT and DPT in Fig. 4(a). As evident from the figure, the $\beta_{eff}^{max}$ and $\tilde{\beta}_{eff}^{max}$ values are strongly dependent on $L_{z}$. The fundamental similarity between TPT and DPT can be seen in the fact that for $L_{z}=2$ and $L_{z}=20$, the numerical values are indeed very similar. However, for intermediate thicknesses, there is a rather significant deviation in between both, mainly showing that TPT starts to approach 3D behavior already for much thinner films, whereas the DPT preserves almost pure 2D behavior until about $L_{z}=10$. This observation displays the exact same tendencies as earlier experimental findings, where ultrathin ferromagnetic films exhibited 2D-like dynamic behavior while their thermodynamic behavior was found to be closer to the 3D universality class [34]. Our simulations thus provide a plausible explanation for the coexistence of different critical behaviors within the same physical sample. It is important to underline again that the presence of a maximum in the effective critical exponent as a function of the control parameter is the hallmark of dimensional crossover. In systems with a well-defined crossover from 3D to 2D behavior, the effective exponent initially exhibits an asymptotic behavior towards 3D values as the system approaches the critical point, reflecting the 3D regime. However, upon approaching ever closer to the true critical point, the finite thickness constrains the correlation length, which leads to a transition towards 2D-like scaling and associated critical exponents. This causes the effective exponent to exhibit a peak as the system transitions to the lower-dimensional scaling regime. Thus, both the value and location of this critical exponent peak encode meaningful information about the crossover process: in our case, the maximum peak value reflects the highest degree of 3D-like behavior observed in the system, and its location indicates how far from the critical point the system begins to be influenced by its reduced dimensionality.
@@
To further elucidate the origin of these quantitative differences in between TPT and DPT in terms of their dimensional crossover behavior, the effective critical exponents $\beta_{eff}$ and $\tilde{\beta}_{eff}$ are displayed as a function of the magnetization $M$ and the dynamic order parameter $Q$ in Figs. 5 (a)-(b) for various values of thickness $L_{z}$ of the film. When comparing these two figures, we observe that the behavior is not only fundamentally and qualitatively similar, but Fig. 5 also illustrates that the dimensional crossover takes place at fairly comparable values of the order parameter when comparing the TPT and DPT. This observation is particularly intriguing given the very significant differences in the reduced distances to the critical point, represented as $\varepsilon^{*}$ and $\tilde{\varepsilon}^{*}$ in Fig. 4 (b). This suggests that the dimensional crossover point for each particular film, identified by $\beta_{eff}^{max}$ and $\tilde{\beta}_{eff}^{max}$, is rather closely correlated to reaching an intermediate level of the respective order parameter, namely about 60 - 80 per cent of its saturation value. Apparently, these values are achieved already at a much greater relative distance to the critical point in the dynamic case, which then explains the reduced ability of DPT systems to exhibit 3D scaling for intermediate films thicknesses $L_{z}$. It is very likely that this difference in quantitative behavior is associated with relevantly different surface behaviors in both cases, which is consistent with earlier findings that demonstrated that the surface behavior in the DPT can deviate significantly from that of the TPT [37, 18].
@@ IV Conclusions
To conclude, this study presents a detailed investigation into the critical behavior and dimensional crossover of DPT in ferromagnetic thin films, using Monte Carlo simulations of the kinetic Ising model. By systematically analyzing the scaling behavior of the dynamic order parameter and comparing it with the TPT, we have illuminated the underlying mechanisms that differentiate these two types of transitions [53, 54].
Our results reveal that while DPT and TPT share commonalities in their scaling laws and critical exponents, the dimensional crossover for the DPT occurs at significantly larger thicknesses than for the TPT. This distinction arises from the relevantly different length-scale and governing principles of the two transitions, with DPT being influenced by time-dependent external fields and non-equilibrium dynamics, whereas TPT adheres to equilibrium-driven processes. The numerical findings reported here align with experimental observations reported in ultrathin Co films [34], where dynamic critical exponents were found to be consistent with two-dimensional scaling, whereas thermodynamic critical exponents appeared to follow three-dimensional behavior. We additionally verified by using a limited set of $L_{z}$ simulations, that log–log slope analyses of the order parameters (not shown here) as a function of system sizes reveal different $L_{z}$ dependencies of the critical exponents for TPT and DPT, which are fully consistent with our main finding in Fig. 4(a). Accordingly, our model study captures a plausible mechanism that would produce such a behavior, namely the delayed onset of 3D scaling for the DPT in intermediately thick magnetic films.
As a final note, it should be emphasized that the observed differences in the crossover behavior and effective critical exponents between TPT and DPT may be influenced by several factors beyond dimensionality alone. Possible contributing elements include the nature of surface effects under equilibrium versus nonequilibrium conditions, and the specific dynamical processes, such as relaxation times and nucleation dynamics, involved in each case. While our present study focuses on characterizing the basic phenomenology of this difference, exploring how these factors quantitatively impact the scaling behavior (particularly whether the exponent difference can be systematically enhanced or reduced) remains an open and intriguing direction for future research.